# Path Planning for Spot Spraying with UAVs Combining TSP and Area Coverages

### v1

@@ Abstract
This paper addresses the following task: given a set of patches or areas of varying sizes that are meant to be serviced within a bounding contour calculate a minimal length path plan for an unmanned aerial vehicle (UAV) such that the path additionally avoids given obstacles areas and does never leave the bounding contour. The application in mind is agricultural spot spraying, where the bounding contour represents the field contour and multiple patches represent multiple weed areas meant to be sprayed. Obstacle areas are ponds or tree islands. The proposed method combines a heuristic solution to a traveling salesman problem (TSP) with optimised area coverage path planning. Two TSP-initialisation and 4 TSP-refinement heuristics as well as two area coverage path planning methods are evaluated on three real-world experiments with three obstacle areas and 15, 19 and 197 patches, respectively. The unsuitability of a Baustrophedon-path for area coverage gap avoidance is discussed and inclusion of a headland path for area coverage is motivated. Two main findings are (i) the particular suitability of one TSP-refinement heuristic, and (ii) the unexpected high contribution of patches areas coverage pathlengths on total pathlength, highlighting the importance of optimised area coverage path planning for spot spraying.
@@ 1 Introduction
The usage of unmanned aerial vehicles (UAVs) is growing in many engineering fields. Both for sensing as well as applying. Likewise, both for civilian (Otto et al. (2018)) as well as military applications (Bortoff (2000), Kabamba et al. (2006)). See Kim et al. (2019) for a review specifically in agriculture.
@@
Within the context of precision agriculture, the motivation and contribution of this paper is to present a path planning method for spot spraying with UAVs: given a set of patches (e.g. weed areas) of varying sizes that are meant to be serviced (with pesticide sprays) within a bounding contour (field contour) calculate a minimal length path plan for an UAV such that the path additionally avoids given obstacles areas (ponds and tree islands endangering the loss of the UAV) and never leaves the bounding contour (for safety reasons).
@@
TSP is ubiquitous in logistics and extensively studied (Bellmore & Nemhauser (1968)). Nevertheless, it is a NP-hard problem and its solution is not obvious for different problem dimensions and applications. See Furchì et al. (2022) for a recent agricultural TSP-application. Likewise, area coverage methods greatly vary for different applications. One recurring method is based on the Baustrophedon-path (Choset & Pignon (1998)), e.g. see Huang et al. (2023), Mukhamediev et al. (2023), Li et al. (2023) within the agricultural context. However, as will be shown, Baustrophedon-based area coverage is unisuitable for precision spot spraying since area coverage gaps typically result. Remedy can be provided by inclusion of a headland path, which is typical for ground-based in-field path planning (Plessen (2019b), Plessen (2018)), and here transfered to the UAV-setting.
In Shah et al. (2020) and Apostolidis et al. (2022) area coverage planning methods different from Baustrophedon-paths are used, but they also suffer from the absence of a headland path for area coverage gap avoidance. In Deng et al. (2019) the operation area is first partitioned into smaller subareas before mitered offset paths are used for area coverage. The disadvantage is that very tight heading changes for the subarea-innermost paths result, which in the limit can only be traced by a holonomic UAV.
A method combining both TSP and area coverage is presented in Xie et al. (2018), where (i) Baustrophedon-based path planning is used for area coverage, (ii) at most 10 separate convexly-shaped patches are considered, and (iii) the TSP is solved by dynamic programming (DP). The authors acknowledge that the DP-approach is not scalable to larger dimensions. In a follow-up paper (Xie et al. (2020)) the same author group extends their method by heuristics to scale for up to 100 patches, however (i) maintaining a Baustrophedon-based area coverage logic, and (ii) assuming only convex polygonal area shapes.
@@ 3.1 Problem Modeling and Subproblem Separation
For a given transition graph connecting all patches and thereby naturally also fixing patch entry and exit points for any path along that transition graph, and under the assumption of (i) permitted straight flight lines between patch entry point and patch exit point, and (ii) area coverage path plans that must include a headland path (i.e., no pure Baustrophedon paths permitted), Problem 1 can without loss of optimality be separated into two subproblems, one solving a TSP for connecting the patches, and one for separately calculating an area coverage path plan for each patch area.
@@
Fig. 4: Illustration of a Baustrophedon path. Characteristic is the zigzag-like path and the absence of a headland path.
Third, arguments why Baustrophedon paths (see Fig. 4) are deemed unsuitable in this paper are given below in Section 3.3 when discussing area coverage path planning in detail.
Under the assumptions and according to Proposition 1 the TSP-solution and patches area coverage are treated as two separate problems. They are discussed in the following two subsections.
@@ 3.2 Subproblem 1: TSP
The basic TSP-problem is: given $N$ locations find a minimum-length tour that visits each location exactly once (Bellmore & Nemhauser (1968)). Finding an optimal TSP-solution is a NP-hard problem. For $N+1$ locations ($N$ patches plus the field entrance location) there are $N\cdot(N-1)\cdot(N-2)\cdot\dots\cdot 1$ unique TSP-paths. In general, the TSP-problem can be cast into a dynamic programming framework (see Bertsekas (2012)) and solved by a label correcting method similarly as in Plessen & Bemporad (2016). However, for Expt. 3 in Fig. 1(c) with 197 patches unimaginable $\mathcal{O}(10^{368})$ unique TSP-combinations result. Thus, for problem dimensions at hand heuristic solution must be devised.
@@ 3.3 Subproblem 2: Patches Area Coverage Path Planning
(a) Classic / Baustrophedon with headland path.
@@
(a) Baustrophedon path: characteristically incurring area coverage gaps. (Type 1) Lanes are circa perpendicular to the area contour. (Type 2) More profound area coverage gaps occur when lanes are almost aligned with the area contour.
@@
Fig. 6: Illustration of area coverage and overlapping area. Notice how for Baustrophedon-based paths significant area coverage gaps occur. This is not acceptable for precision spot spraying. (As extending detail, it is here assumed that UAV-spraying is always switched on, which is indicated by the darker blue overlapping areas. In practice, overlap areas can be reduced by appropriate boom section control and individual spray nozzle control.)
Fig. 7: Baustrophedon path for Fig. 6(a). Baustrophedon-pathlengths are much shorter since not including any headland coverage. However, on the flip-side they do not offer full area coverage. Since this is of predominant importance for agricultural applications, such as pesticide spraying and the like, Baustrophedon paths are not considered for the remainder of this paper.
@@
The by far pre-dominant area coverage path planning method for UAVs is based on Baustrophedon paths (sketched in Fig. 4), see e.g. Xie et al. (2020), Huang et al. (2023), Li et al. (2023), Huang et al. (2023) and Coombes et al. (2017).
In this paper, however, it is argued that Baustrophedon-based coverage path planning is unsuitable for precision spot spraying applications. Two reasons are given. First, characteristic for Baustrophedon-paths are area coverage gaps, see Fig. 6(a). Second, one method to counteract these gaps is to enlarge obstacles area (e.g. in Pham et al. (2020)) or, in spirit similar, artificially enlarge patches areas such that original core patches areas are fully covered. This method however fails when patches areas are directly bordering the field area contour (which is exactly the case for the experimental real-world data in this paper). Furthermore, this method of artificially enlarging patches areas then entails the danger of overspraying healthy field areas (i.e. beyond the original core patches areas).
@@
    A Baustrophedon-path extended with a headland path shall in the following be referred to as the Classic area coverage path planning method, since it is the classic method used for ground-based area coverage in agricultural fields.
@@
The benefit of including a headland path for area coverage path planning in comparison to Baustrophedon-based planning is visualised in Fig. 6.
@@ 3.4 Obstacle avoidance method
4 Numerical Results and Extending Discussion
@@ 4.1 Extending Discussion
This subsection discusses details and extending variations. First, when zooming into Fig. 9(c) one can observe two locations where planned paths unexpectedly exceed field area contours (in the vicinity of patches numbered 179 and 185). When further analysing this it was found that the reason was ’faulty’ input data, where coordinates for the two patches were slightly exceeding field contours and the calculated projection points connecting patches in the vicinity were consequently located slightly outside the field area. After manual correction of the’faulty’ data (by slightly modifying field contour coordinates at the two locations), the resulting UAV-path stays within field contours as Fig. 11 illustrates (thereby also validating the method from Sect. 3.4 used to correct TSP-paths to stay within field area contours). This anomaly detection example underlines the importance of input data verification and potential dangerous impacts (UAV-paths exceeding the field area contour) inaccurate input data may have.
@@
Third, a perceived strength of proposed method is that only two parameters are needed, see Table 1. This os complexified when accounting for more detailed UAV-dynamics such as a specific turning radius or other dynamical vehicle constraints. Here, the method from Plessen (2024) (originally devised for ground-based motion) can be employed before adding a z-coordinate reference, either (i) by a constant offset distance tracking the field topograhpy, or (ii) by a piecewise-affine reference as suggested in Lin & Saripalli (2014), where a 2D Dubins curve is first created before a z-coordinate reference is then added. The inclusion of UAV-dynamics with limited turning radii, which is especially relevant for fixed-wing aircraft and useful for larger patches areas, renders path planning more complex. See Fig. 13 for exemplary illustration of a path plan when the UAV-turning radius is larger than half the operating width. Several comments are made. First, this figure illustrates the benefits of quadrotor-like UAVs with near-holonomic flight dynamics that can avoid large turning radii characteristic for fixed-wing UAVs. Second, flight maneuvers exceeding the patch area as shown in Fig. 13 offer the possibility of counteracting the downsides of Baustrophedon paths by enabling full area coverage by shifting the turning outside of the patch area. The flip-side, however, is that such maneuvers then might violate (i.e. exceed) the field contour, especially when a patch area is located close to the field contour (as is precisely the case for this particular patch, see patch 18 in Fig. 9(b)). 3D precision trajectory generation for UAVs is a topic of future work, likely also involving the planning of multiple cascaded headland paths such that flight maneuevers are contained to the patches areas even for turning radii larger than half the UAV operating width. Quadrotors offer the benefit of near-holonomic maneuverability. But fixed-wing UAVs might be more suited for large-scale areas.

### v2

@@ Abstract
This paper addresses the following task: given a set of patches or areas of varying sizes that are meant to be serviced within a bounding contour calculate a minimal length path plan for an unmanned aerial vehicle (UAV) such that the path additionally avoids given obstacles areas and does never leave the bounding contour. The application in mind is agricultural spot spraying, where the bounding contour represents the field contour and multiple patches represent multiple weed areas meant to be sprayed. Obstacle areas are ponds or tree islands. The proposed method combines a heuristic solution to a traveling salesman problem (TSP) with optimised area coverage path planning. Two TSP-initialisation and 4 TSP-refinement heuristics as well as two area coverage path planning methods are evaluated on three real-world experiments with three obstacle areas and 15, 19 and 197 patches, respectively. The unsuitability of a Boustrophedon-path for area coverage gap avoidance is discussed and inclusion of a headland path for area coverage is motivated. Two main findings are (i) the particular suitability of one TSP-refinement heuristic, and (ii) the unexpected high contribution of patches areas coverage pathlengths on total pathlength, highlighting the importance of optimised area coverage path planning for spot spraying.
@@ 1 Introduction
Data InputSubprob. 1: TSP-path connecting patches areas For Each Patch:Subprob. 2: Patch Area Coverage Path Connect one TSP-path and all coverage paths Return path coordinates
Fig. 1: High-level algorithm of this paper as a block diagram. A variety of heuristics are compared for Subproblem 1. Two main approaches are compared for Subproblem 2. Data input comprises one field contour, multiple patches areas contours, potential obstacles areas contours coordinates as well as a working width of the UAV.
@@
The usage of unmanned aerial vehicles (UAVs) is growing in many engineering fields. Both for sensing as well as applying. Likewise, both for civilian (Otto et al. (2018)) as well as other applications (Bortoff (2000), Kabamba et al. (2006)). See Kim et al. (2019) for a review specifically in agriculture.
@@
Within the context of precision agriculture, the motivation and contribution of this paper is to present a path planning method for spot spraying with UAVs: given a set of patches (e.g. weed or damaged areas, Plessen (2025)) of varying sizes that are meant to be serviced (with pesticide sprays) within a bounding contour (field contour) calculate a minimal length path plan for an UAV such that the path additionally avoids given obstacles areas (ponds and tree islands endangering the loss of the UAV) and never leaves the bounding contour (for safety reasons).
@@
TSP is ubiquitous in logistics and extensively studied (Bellmore & Nemhauser (1968)). Nevertheless, it is a NP-hard problem and its solution is not obvious for different problem dimensions and applications. See Furchì et al. (2022) for a recent agricultural TSP-application. Likewise, area coverage methods greatly vary for different applications. One recurring method is based on the Boustrophedon-path (Choset & Pignon (1998)), e.g. see Huang et al. (2023), Mukhamediev et al. (2023), Li et al. (2023) within the agricultural context. However, as will be shown, Boustrophedon-based area coverage is unisuitable for precision spot spraying since area coverage gaps typically result. Remedy can be provided by inclusion of a headland path, which is typical for ground-based in-field path planning (Plessen (2019b), Plessen (2018)), and here transfered to the UAV-setting.
In Shah et al. (2020) and Apostolidis et al. (2022) area coverage planning methods different from Boustrophedon-paths are used, but they also suffer from the absence of a headland path for area coverage gap avoidance. In Deng et al. (2019) the operation area is first partitioned into smaller subareas before mitered offset paths are used for area coverage. The disadvantage is that very tight heading changes for the subarea-innermost paths result, which in the limit can only be traced by a holonomic UAV.
A method combining both TSP and area coverage is presented in Xie et al. (2018), where (i) Boustrophedon-based path planning is used for area coverage, (ii) at most 10 separate convexly-shaped patches are considered, and (iii) the TSP is solved by dynamic programming (DP). The authors acknowledge that the DP-approach is not scalable to larger dimensions. In a follow-up paper (Xie et al. (2020)) the same author group extends their method by heuristics to scale for up to 100 patches, however (i) maintaining a Boustrophedon-based area coverage logic, and (ii) assuming only convex polygonal area shapes.
@@ 3.1 Problem Modeling and Subproblem Separation
For a given transition graph connecting all patches and thereby naturally also fixing patch entry and exit points for any path along that transition graph, and under the assumption of (i) permitted straight flight lines between patch entry point and patch exit point, and (ii) area coverage path plans that must include a headland path (i.e., no pure Boustrophedon paths permitted), Problem 1 can without loss of optimality be separated into two subproblems, one solving a TSP for connecting the patches, and one for separately calculating an area coverage path plan for each patch area.
@@
Fig. 5: Illustration of a Boustrophedon path. Characteristic is the zigzag-like path and the absence of a headland path.
Third, arguments why Boustrophedon paths (see Fig. 5) are deemed unsuitable in this paper are given below in Section 3.3 when discussing area coverage path planning in detail.
Under the assumptions and according to Proposition 1 the TSP-solution and patches area coverage are treated as two separate problems within the framework of Fig. 1. The two subproblems are discussed in the following two subsections.
@@ 3.2 Subproblem 1: TSP
The basic TSP-problem is: given $N$ locations find a minimum-length tour that visits each location exactly once (Bellmore & Nemhauser (1968)). Finding an optimal TSP-solution is a NP-hard problem. For $N+1$ locations ($N$ patches plus the field entrance location) there are $N\cdot(N-1)\cdot(N-2)\cdot\dots\cdot 1$ unique TSP-paths. In general, the TSP-problem can be cast into a dynamic programming framework (see Bertsekas (2012)) and solved by a label correcting method similarly as in Plessen & Bemporad (2016). However, for Expt. 3 in Fig. 2(c) with 197 patches unimaginable $\mathcal{O}(10^{368})$ unique TSP-combinations result. Thus, for problem dimensions at hand a heuristic solution must be devised.
@@ 3.3 Subproblem 2: Patches Area Coverage Path Planning
(a) Classic / Boustrophedon with headland path.
@@
(a) Boustrophedon path: characteristically incurring area coverage gaps. (Type 1) Lanes are circa perpendicular to the area contour. (Type 2) More profound area coverage gaps occur when lanes are almost aligned with the area contour.
@@
Fig. 7: Illustration of area coverage and overlapping area. Notice how for Boustrophedon-based paths significant area coverage gaps occur. This is not acceptable for precision spot spraying. (As extending detail, it is here assumed that UAV-spraying is always switched on, which is indicated by the darker blue overlapping areas. In practice, overlap areas can be reduced by appropriate boom section control and individual spray nozzle control.)
Fig. 8: Boustrophedon path for Fig. 7(a). Boustrophedon-pathlengths are much shorter since not including any headland coverage. However, on the flip-side they do not offer full area coverage. Since this is of predominant importance for agricultural applications, such as pesticide spraying and the like, Boustrophedon paths are not considered for the remainder of this paper.
@@
The by far pre-dominant area coverage path planning method for UAVs is based on Boustrophedon paths (sketched in Fig. 5), see e.g. Xie et al. (2020), Huang et al. (2023), Li et al. (2023), Huang et al. (2023) and Coombes et al. (2017).
In this paper, however, it is argued that Boustrophedon-based coverage path planning is unsuitable for precision spot spraying applications. Two reasons are given. First, characteristic for Boustrophedon-paths are area coverage gaps, see Fig. 7(a). Second, one method to counteract these gaps is to enlarge obstacles area (e.g. in Pham et al. (2020)) or, in spirit similar, artificially enlarge patches areas such that original core patches areas are fully covered. This method however fails when patches areas are directly bordering the field area contour (which is exactly the case for the experimental real-world data in this paper). Furthermore, this method of artificially enlarging patches areas then entails the danger of overspraying healthy field areas (i.e. beyond the original core patches areas).
@@
    A Boustrophedon-path extended with a headland path shall in the following be referred to as the Classic area coverage path planning method, since it is the classic method used for ground-based area coverage in agricultural fields.
@@
The benefit of including a headland path for area coverage path planning in comparison to Boustrophedon-based planning is visualised in Fig. 7.
@@ 3.4 Obstacle avoidance method
4 Numerical Results, Discussion and Limitations
@@ 4.1 Extending Discussion and Limitations
This subsection discusses details, extending variations and limitations. First, when zooming into Fig. 10(c) one can observe two locations where planned paths unexpectedly exceed field area contours (in the vicinity of patches numbered 179 and 185). When further analysing this it was found that the reason was faulty input data (data was used ’as is’), where coordinates for the two patches were slightly exceeding field contours and the calculated projection points connecting patches in the vicinity were consequently located slightly outside the field area. After manual correction of the data (by slightly modifying field contour coordinates at the two locations), the resulting UAV-path stays within field contours as Fig. 12 illustrates (thereby also validating the method from Sect. 3.4 used to correct TSP-paths to stay within field area contours). This anomaly detection example underlines the importance of input data verification and potential dangerous impacts (UAV-paths exceeding the field area contour) inaccurate input data may have.
@@
Third, a perceived strength of proposed method is that only two parameters are needed, see Table 1. This is complexified when accounting for more detailed UAV-dynamics such as a specific turning radius or other dynamical vehicle constraints. Here, the method from Plessen (2024) (originally devised for ground-based motion) can be employed before adding a z-coordinate reference, either (i) by a constant offset distance tracking the field topograhpy, or (ii) by a piecewise-affine reference as suggested in Lin & Saripalli (2014), where a 2D Dubins curve is first created before a z-coordinate reference is added.
An important aspect with respect to z-coordinate and 3D field topography is made. The method presented in this paper is a graph optimisation method, based on a transition graph model according to Section 3.1. Therefore, on the graph optimisation level the presented method does not change for both 2D- and 3D-path planning. What changes, however, in those two cases is the mapping from graph model to actual path coordinates, which then are either 2D or 3D coordinates.
The inclusion of UAV-dynamics with limited turning radii, which is especially relevant for fixed-wing aircraft and useful for larger patches areas, renders path planning more complex. See Fig. 14 for exemplary illustration of a path plan when the UAV-turning radius is larger than half the operating width. This figure illustrates the benefits of quadrotor-like UAVs with near-holonomic flight dynamics that can avoid large turning radii characteristic for fixed-wing UAVs. Furthermore, flight maneuvers exceeding the patch area as shown in Fig. 14 offer the possibility of counteracting the downsides of Boustrophedon paths by enabling full area coverage by shifting the turning outside of the patch area. The flip-side, however, is that such maneuvers then might violate (i.e. exceed) the field contour, especially when a patch area is located close to the field contour (as is precisely the case for this particular patch, see patch 18 in Fig. 10(b)). 3D precision trajectory generation for UAVs is a topic of future work, also involving the planning of multiple cascaded headland paths such that flight maneuevers are contained to the patches areas even for turning radii larger than half the UAV operating width. Quadrotors offer the benefit of near-holonomic maneuverability. But fixed-wing UAVs might be more suited for large-scale areas.
A limitation of this paper is the absence of´ real-world flight experiments. The scope of this paper is nominal mission planning, in particular, path planning (not path tracking) according to Problem 1, for which a high-level algorithm according to Fig. 1 is presented. The method is evaluated on real-world data.